estion 2 of 3 > How many atoms are in 4.66 mol of pure aluminum? Al atoms
Answer #1

Answer:-

Given:-

No. of moles of pure aluminium = 4.66 moles

As we know that a mole is defined as the amount of substance that contains the same no. of particles as the number of atoms in exactly 12 g of C-12. A mole is a basic unit of the amount or quantity of the substance that may be atoms , molecules , ions or groups of ions (such as in salt). Its also be defined as the amount of a substance containing elementary particles such as atoms , molecules , ions or groups of ions (such as in salt) equal to the Avogadro's number i.e 6.023 \times 1023 . e.g.

1 mole Na = 23 g Na = 6.023 \times 1023 Na atoms

1 mole O2 = 32 g O2 = 6.023 \times 1023 O2 molecules

1 mole Cl- = 35.5 g Cl- = 6.023 \times 1023 Cl- ions

therefore

Also we know that

1 mole pure aluminium (Al) = 27 g Al contained = 6.023 \times 1023 Al atoms

So

1 mole pure aluminium (Al) contained = 6.023 \times 1023 Al atoms

then

4.66 moles pure aluminium (Al) contained = 4.66 \times 6.023 \times 1023 Al atoms

4.66 moles pure aluminium (Al) contained = 28.0672 \times 1023 Al atoms (i.e the answer)